(7%) Problem S: A hollow conducting sphere has an inner radius of r,-0.11 m and an outside radius ofr,-0.27 m. The sphere has a net charge of Q = 4.9E-06 C. - 50% Part (a) what is the field El in NC 1 m from the spheres outside surface? 50% Part (b) What is the magnitude of the field at a distance of 0.05 m from the center of the sphere? Grade Summary Deductions Potential 090 100%
Answer #1

a)The entire charge Q settles on the outside surface of the sphere. From Gauss's law it can be proved that the field due to the conducting sphere for all outside points can be obtained by assuming that all the charge is concentrated within the sphere at its center.

Now E, = Q/4πεο r2ー4.9 * 10-6C * 9 * 109N㎡/C2/(1+0.27)In-4.9 * 10-6 * 9 * 109/1.27= 3.5 * 104M/C

since r=0.27+1 m=1.27 m.

b) Within the sphere the electric field is zero since it is conducting and all the charge settles on it i.e on its surface.